OpenStudy (anonymous):

help...what is x 9-9x=20x-15

OpenStudy (magbak):

x= 24/29

OpenStudy (foreverandalways13):

do you understand how they got there?

OpenStudy (magbak):

I do

OpenStudy (foreverandalways13):

lol

OpenStudy (foreverandalways13):

@nocontact110

OpenStudy (anonymous):

yes

OpenStudy (magbak):

See he does

OpenStudy (magbak):

Ok you need to isolaate x so you add 9x to both sides making it 29x-15=9 Now you add 15 to both sides making it 29x=24 Now divide both sides by 29 and this gives you x=24/29

OpenStudy (foreverandalways13):

yup!

OpenStudy (anonymous):

wow ok i see where i messed up thanks

OpenStudy (magbak):

Your welcom no problem at all alwayse happy to help.